OpenStudy (anonymous):

[8.01] Factor completely: −7x2 + 14x − 21 −1(x2 − 14x + 21) −7x(x2 − 2x + 3) −7(x2 + 2x − 3) −7(x2 − 2x + 3) I think its A help please

OpenStudy (anonymous):

each term has a common factor of \(-7\)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so not A

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

gotta be \[-7(\text{something})\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what would give you \(14x\) ? i .e. fill in the blank \(-7\times \)___ \(=14x\)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

no actually \(-7\times2x=-14x\) but you want \(14x\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you need \(-2x\) because \(-7\times -2x=14x\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and finally you need \(-7\times \)____\(=-21\)

OpenStudy (anonymous):

what would go in the blank?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

yes \(+3\)

OpenStudy (anonymous):

so \[−7x^2 + 14x − 21=-7(x^2-2x+3)\]

OpenStudy (anonymous):

so its D

OpenStudy (anonymous):

yes it is

OpenStudy (texaschic101):

nicely done :)